OpenStudy (anonymous):

find dy/dx of the following 1) x^2+y^2-a^2=0 2) x^2/a^2+y^2/b^2=1 3) x^3-4xy^2-2y^2-5y^3=0 plz help...:(

OpenStudy (anonymous):

.1. 2x + 2y.dy/dx = 0 dy/dx = -x/y 2. 2x/a^2 +( 2y/b^2 ). dy/dx = 0 dy/dx = - 2x/a^2 / 2y/b^2 = -2x * b^2 --- --- a^2 2y = - x b^2 ---- y a^2

OpenStudy (anonymous):

3. 3x^2 - (4x 2y. dy/dx +4y^2) - 4y .dy/dx -15y^2 . dy/dx = 0 3x^2 - 8xy. dy/dx - 4y^2 - 4y .dy/dx -15y^2 . dy/dx = 0 8xy. dy/dx + 4y .dy/dx + 15y^2 . dy/dx = 3x^2 dy/dx = 3x^2 / (8xy + 4 + 15y^2)

OpenStudy (anonymous):

which bit are u talking about?

OpenStudy (anonymous):

???

OpenStudy (anonymous):

oh i understand now thanx...:)

OpenStudy (anonymous):

y w